OPTICALLY PUMPED FAR-INFRARED LASERS FROM METHANOL ISOTOPOMERS: 12CD_{3}OH, ^{12} CH_{3}OD, AND ^{12}CH_{2}DOH; NEW LASER LINES AND FREQUENCY MEASUREMENTS

\underline{E.C.C. VASCONCELLOS}, S.C. ZERBETTO , Instituto de Fisica Gleb Wataghin, Departamento de Eletronica Quantica, Universidade Estadual de Campinas (UNICAMP), 13083-970, Campinas, SP, Brazil ; L.R. ZINK, and K.M. EVENSON, National Institute of Standards and Technology, Time and Frequency Division, Boulder, CO 80303-3328.

Methanol and ten of its isotopomers have generated over 2000 far-infrared laser lines when pumped by CO_{2} lasers. We report here 28 new laser lines from the isotopomers^{12}CD_{3}OH,^{12}CH_{3}OD,^{12}CH_{2}DOH along with their frequencies, relative intensities, relative polarizations and pump offsets. An efficient CO_{2} laser with extended line coverage provided pumping from high J-fundamental, sequence and hot-band CO_{2} transitions. Many new short-wavelength (below 100 μm) FIR laser transitions were observed.